On March 10, 2022, Crane Company sells equipment that it purchased for $193,920 on August 20, 2015. It was originally estimated that the equipment would have a life of 12 years and a salvage value of $16,968 at the end of that time, and depreciation has been computed on that basis. The company uses the straight-line method of depreciation. Compute the depreciation charge on this equipment for 2015, for 2022, and the total charge for the period from 2016 to 2021, inclusive, under each of the six following assumptions with respect to partial periods. (Round depreciation per day to 2 decimal places, e.g. 15.64 and final answers to 0 decimal places, e.g. 45,892.) 2015 2016-2021 Inclusive 2022 1. Depreciation is computed for the exact period of time during which the asset is owned. (Use 365 days for the base and record depreciation through March 9, 2022.) $ 14746 $ 73730 $ 8120.4 2. Depreciation is computed for the full year on the January 1 balance in the asset account. $ $ $ 3. Depreciation is computed for the full year on the December 31 balance in the asset account. $ $ $ 4. Depreciation for the one-half year is charged on plant assets acquired or disposed of during the year. $ $ $ 5. Depreciation is computed on additions from the beginning of the month following the acquisition and on disposals to the beginning of the month following disposal. $ $ $ 6. Depreciation is computed for a full period on all assets in use for over one-half year, and no depreciation is charged on assets in use for less than one-half year. (Use 365 days for a base.) $
Answer #1

Depreciation p.a.=(value of asset - scrap value) ÷ life of asset

=(193920-16968)÷12

=14746

Depreciation for 2015=14746×122÷365=4929

Depreciation for 2022=14746×90÷365=3636

Depreciation from 2016 to 2021 will be14746×6=88476